OpenStudy (anonymous):

FAN AND MEDAL!!! Find the quadratic function y=a(x-h)^2 whos graph passes through the given points. (8,0) and (12,-14)

OpenStudy (anonymous):

Plug in the points for x and y, then you'll be able to solve for a and h. What do you get when you substitute (8, 0)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Plug in 8 for x and 0 for y. Then show me what you have left.

OpenStudy (anonymous):

0=a(8-h)^2

OpenStudy (anonymous):

right. now you can divide both sides by a to get 0 = (8 - h)² and you can solve that for h

OpenStudy (anonymous):

but what did you do with the a?

OpenStudy (anonymous):

divided both sides by a. 0/a = 0

OpenStudy (anonymous):

the (8,0) is the vertex of the parabola

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so just plug in the 2nd point \[-14 = a(12 - 8)^2\]